OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 2y = 2 x + 3y = 17

OpenStudy (nubeer):

re- arrange your first equation x = 2+ 2y ( got this by adding 2y on both sides.) now put this value of x in the second equation.

OpenStudy (zpupster):

in the first equation add 2y to both sides of equation like so: x-2y +2y = 2+2y we are trying to get x by itself. we will have x=2y+2 substitute this for the x in 2nd equation

OpenStudy (nubeer):

you will get 2+2y +3y =17 now solve it for y

OpenStudy (anonymous):

yes i got it right when you put it up! haha thanks(: